OpenStudy (anonymous):

Write a polynomial function of minimum degree with real coefficients whose zeros include those listed. Write the polynomial in standard form. 4, -8, and 2 + 3i

OpenStudy (anonymous):

i know i have to multiply (x-4)(x+8) but i dont know what to do about 2+3i

ganeshie8 (ganeshie8):

if 2+3i is a root, 2-3i is also a root cuz complex roots never come single, they always come in conjugate pairs

ganeshie8 (ganeshie8):

so you need to multiply below factors also :- (x-(2+3i)) and (x-(2-3i))

OpenStudy (anonymous):

how do u multiply i's together?

ganeshie8 (ganeshie8):

okay, so the required polynomial wud be :- (x-4)(x+8)(x-(2+3i))(x-(2-3i))

ganeshie8 (ganeshie8):

you knw how to deal with first two factors, but u r not sure how to multiply last two factors is it ?

OpenStudy (anonymous):

yes

ganeshie8 (ganeshie8):

(x-(2+3i))(x-(2-3i)) (x-2-3i)(x-2+3i)

ganeshie8 (ganeshie8):

(x-2-3i)(x-2+3i) --- ----

ganeshie8 (ganeshie8):

see that it is of the form (a-bi)(a+bi) a = x-2 b = 3

OpenStudy (anonymous):

is that a formula?

ganeshie8 (ganeshie8):

then, simply use this formula : (a-bi)(a+bi) = a^2+b^2

ganeshie8 (ganeshie8):

(x-(2+3i))(x-(2-3i)) (x-2-3i)(x-2+3i) ((x-2)^2 + 3^2)

OpenStudy (anonymous):

so what is the a^2+b^2 for

ganeshie8 (ganeshie8):

thats just a formula,

ganeshie8 (ganeshie8):

(x-(2+3i))(x-(2-3i)) (x-2-3i)(x-2+3i) ((x-2)^2 + 3^2) (x^2 -4x+4+9)

ganeshie8 (ganeshie8):

simplify

OpenStudy (anonymous):

oh so thats how you multiply them together?

ganeshie8 (ganeshie8):

yes ! remember this formula :- (a+bi)(a-bi) = a^2 + b^2

ganeshie8 (ganeshie8):

and dont forget to multiply this with the first two factors good luck !

OpenStudy (anonymous):

i got f(x) = x^4 - 35x^2 + 180x - 416

ganeshie8 (ganeshie8):

Excellent !! wolfram says the same http://www.wolframalpha.com/input/?i=%28x-%282%2B3i%29%29%28x-%282-3i%29%29%28x-4%29%28x%2B8%29+

OpenStudy (anonymous):

thanks

ganeshie8 (ganeshie8):

np :)